OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = –5 and a6 = –5,120?

OpenStudy (amistre64):

d = (-5120 + 5)/5 wince we are looking for 4 arithmetic means between a1 and a6

OpenStudy (amistre64):

an = a1 + d(n-1) is the general formula to find any nth term in an arith sequence

OpenStudy (amistre64):

a9 = a1 + d(8) a9 = -5 + (-5120+5)8/5

OpenStudy (precal):

@amistre64 the original question is asking for a geometric sequence, just fyi

OpenStudy (precal):

 use this formula

OpenStudy (precal):

you will need a common ratio

OpenStudy (precal):